OpenStudy (anonymous):

PLEASE HELP. x^4−x^3−13x^2+x+12. solve for all the real zeros.

OpenStudy (anonymous):

@jim_thompson5910

OpenStudy (anonymous):

if you don't want to cheat, try the obvious possibilities first. for example try \(f(1)\) which is easiest

OpenStudy (anonymous):

thanks its fine

OpenStudy (anonymous):

\[f((x)= x^4−x^3−13x^2+x+12\] \[f(1)=1-1-13+1+12=0\] works, which gives you one zero right off the bat